# Peer-Aware Cost Estimation in Nonlinear General-Sum Dynamic Games for Mutual Learning and Intent Inference

### v1

@@ Abstract
Human-robot interactions can be modeled as incomplete-information general-sum dynamic games since the objective functions of both agents are not explicitly known to each other. However, solving for equilibrium policies for such games presents a major challenge, especially if the games involve nonlinear underlying dynamics. To simplify the problem, existing work often assumes that one agent is an expert with complete information about its peer, which can lead to biased estimates and failures in coordination. To address this challenge, we propose a nonlinear peer-aware cost estimation (N-PACE) algorithm for general-sum dynamic games. In N-PACE, using iterative linear quadratic (LQ) approximation of the nonlinear general-sum game, each agent explicitly models the learning dynamics of its peer agent while inferring their objective functions, leading to unbiased fast learning in inferring the unknown objective function of the peer agent, which is critical for task completion and safety assurance. Additionally, we demonstrate how N-PACE enables intent communication in such multi-agent systems by explicitly modeling the peer’s learning dynamics.
I INTRODUCTION
@@
General-sum dynamic games offer a powerful framework for modeling multi-agent interactions [1, 2], with applications in autonomous driving [3], assistive robots [4], and social navigation [5]. However, the process of finding the feedback Nash equilibrium (FBNE) solution of such games, which can be achieved by solving a set of Hamilton-Jacobi-Bellman (HJB) equations [6], is inherently complex, as traditional dynamic programming approaches suffer from the curse of dimensionality [7]. This problem becomes even more challenging when agents do not know each other’s objective function and must perform inference online while making decisions, resulting in an incomplete-information differential game, a common scenario in human-robot interaction (HRI) [4, 8, 9, 10] and autonomous vehicle interaction [11, 12, 13].
@@
Figure 1: Demonstration of the core idea in N-PACE using a simulated general-sum game between two autonomous cars at an intersection. In (a), each agent assumes its peer knows its aggressiveness level, leading to biased inferences and possible collision. In (b), N-PACE allows agents to learn their peer’s aggressiveness while accounting for mutual learning, ensuring safe intersection crossing.
@@
Existing approaches to solve such games often rely on the assumption that one agent is an expert which knows the true intent of other agents [14, 15, 1]. Despite some empirical success of these approaches, in scenarios where all agents are learning simultaneously and inferring each other’s objective functions (which is defined as their intent in this paper), these assumptions can lead to biased estimates, poor coordination, and failure to achieve each agent’s goal [16]. This is because while doing the inference, they fail to account for the fact that other agents can be non-stationary learners, whose policies are conditioned on previously observed actions [9]. It is even more important to consider such mutual learning in a competitive setting where an agent has no information of the others’ true intent.
@@
In this paper, we highlight the importance of modeling the learning dynamics of peer agents while inferring their objectives, addressing critical limitations in existing approaches. We achieve this goal by developing a peer-aware cost estimation algorithm for general-sum nonlinear games (N-PACE), where each agent leverages prior knowledge about its peer agent’s learning dynamics (e.g., gradient-based learner and Bayesian learner) to condition its current inference of the other’s objective on their parameter learning, thereby removing the bias in their inferences. The N-PACE algorithm employs iterative linear quadratic (iLQ) approximation of the nonlinear general-sum game [5], both in policy update and intent inference phases. We also demonstrate how (iLQ) enables efficient approximation during the inference phase of our algorithm.
@@
Contributions. This paper makes the following contributions. (1) We introduce the N-PACE algorithm for mutual intent inference in multi-agent interaction where all agents are making inference online. (2) We empirically demonstrate that N-PACE is critical for task completion and safety in multi-agent systems where all agents are learners, compared to the incorrect assumption that a peer agent is an expert who knows the ego agent’s true intent. (3) We demonstrate how N-PACE, along with access to the learning dynamics of the peer agent, can be leveraged for intent communication and signaling to facilitate mutual learning and adaptation.
@@
The closest works to ours are [5, 15, 16], where [5] uses iLQ for solving complete information multi-agent interactions. In [15], the iLQ approach is employed in a general-sum game where an “expert” demonstrates its intent to a learner agent knowing their learning dynamics, but it does not consider the situation when two learning agents (no experts) try to coordinate which is a more practical setting in multi-agent interactions. Finally, while [16] studies the problem of mutual learning and adaptation, its problem setup is limited to a single-step optimization in the cost function. This work extends it to a finite-horizon optimization problem which is more common in multi-agent interactions.
II Related Work
@@
Solving general-sum games. Multi-agent interactions have been explored using various frameworks, including multi-agent reinforcement learning [17, 18], adaptive control [19], and game theory [20, 21]. A large portion of these interactions can be modeled as general-sum dynamic games [22, 23]. Traditional approaches for solving such games rely on adaptive dynamic programming to compute feedback Nash equilibria (FBNE) [24]. However, these methods are often hindered by the curse of dimensionality [7], making them impractical for high-dimensional problems and real-time implementation in robotics. To overcome these limitations, physics informed neural networks (PINN) have been used to solve the (HJB) equation [23] more efficiently. Some recent work has focused on iterative linear-quadratic (iLQ) approximations, which offer a more computationally efficient strategy for for approximating the multi-agent interactions [5] solution by finding the local Nash equilibrium of such games. In this paper, we employ iLQR for both the learning and control phases to find the approximated Nash equilibrium policies.
@@
Incomplete information general-sum games. Solving incomplete information general-sum games is significantly more challenging than their complete-information counterparts [25, 26]. These games often require simplifications, such as modeling one agent as a learner while assuming others are fully informed [14, 27], effectively simplifying them as optimal or adaptive control problems [4]. Other approaches discretize the intent parameter and rely on offline datasets for intent inference and belief updates [28, 3]. Our work is motivated by the lack of studies on incomplete information dynamic games where all agents are learners [29]. We highlight the potential issues caused by these simplifications, such as assuming other agents are fully informed.
@@
Learning from a learner. Online inference of an expert agent’s objective function has been extensively studied through inverse optimal control and inverse reinforcement learning [30, 31, 32]. While inverse non-cooperative games have been explored both offline and online [33, 34], they also primarily focus on learning from an expert system. Some works have considered learning from a learner agent by incorporating their learning dynamics [35, 29], but they do not directly address interactions where multiple learner agents simultaneously infer each other’s cost or reward function parameters. Building on these learning from a learner approaches and recent efforts in modeling human learning dynamics for improved intent inference in human-robot interactions [10, 9], we propose a framework for mutual intent inference in systems where all agents are learners.
III Methods
In this section, we first introduce the notations and formulate the multi-agent interaction problem as a general-sum incomplete information differential game. We then discuss how iLQR can be used for solving such games. Next, we propose a framework for modeling each agent’s learning dynamics. Finally, we present our algorithm, N-PACE, and demonstrate its role in intent communication.
III-A Problem Formulation
@@
Notations and Assumptions. Without loss of generality to more agents, we focus on a two-player setting with agents $i$ and $j$. When referring to an arbitrary agent, we use the index $k$ and denote their counterpart as $-k$. We consider the following system dynamics
@@
Feedback Nash Equilibrium (FBNE) [36] is a solution concept for the above minimization problem, constrained by (1), where neither agent has an incentive to deviate from their policy, as doing so would increase their cumulative cost. In an FBNE, each agent’s policy is a function of the current state, ensuring that at every time step, the chosen control actions are optimal given the strategies of the other agents. We represent these policies using $\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$ such that $a^{k}_{t}=\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$, emphasizing that they depend on both $\theta^{i}$ and $\theta^{j}$, which is crucial for further formulation in our work.
@@
The incomplete information game. The previously discussed FBNE solution was derived under the assumption that both agents have full knowledge of $\theta^{i}$ and $\theta^{j}$. However, in incomplete information games, $\theta^{j}$ may be unknown to agent $i$, and $\theta^{i}$ may be unknown to agent $j$. As a result, at each time step, they must rely on estimations $\hat{\theta}^{j}_{t}$ and $\hat{\theta}^{i}_{t}$, respectively, and update these estimates over time. This leads to policies of the form $\pi^{i}_{t}(s_{t};\theta^{i},\hat{\theta}^{j}_{t})$ and $\pi^{j}_{t}(s_{t};\hat{\theta}^{i}_{t},\theta^{j})$, which are not necessarily the same as FBNE strategies. To address this problem, the next section introduces how we use the iLQR method as a solver to obtain the policy $\pi^{k}_{t}(s_{t};\theta^{k},\hat{\theta}^{-k}_{t})$ for each agent $k$, based on their current estimate of the other agent’s parameter $\hat{\theta}^{-k}_{t}$. We then propose appropriate update rules for these estimates in Sec. III-C.
III-B iLQR for Multi-Agent Interactions
@@
The iterative linear quadratic regulator (iLQR) approach for locally optimal feedback control of nonlinear systems was first introduced in [37]. More recent work, such as [5], has demonstrated its efficiency in computing local Nash equilibrium solutions for general-sum dynamic games in real time. In our incomplete information game, where we need policies $\pi^{k}_{t}$ to be computed as functions of the estimated parameters $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$ in real-time, we leverage the proposed iLQR algorithm in [5] to solve the game for each agent $k$ at every time step. The main reasons for using iLQR in this work are its efficiency in real-time computations and, more importantly, its capability to parameterize the obtained policies as functions of $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$, which is crucial for N-PACE as shown in Sec (III-D).
@@
In the iLQR approach, we begin with an initial admissible policy for each agent. In the forward pass, by applying these policies to the system dynamics in (1), we obtain a trajectory $\boldsymbol{\zeta^{0}_{t}}=\{(s^{0}_{\tau},a^{i0}_{\tau},a^{j0}_{\tau})\mid\tau\in\{t,t+1,\dots,T\}\}$. Next, by linearizing the dynamics $f_{t}$ and quadratizing the running costs $g^{k}_{t}$ in the neighborhood of the trajectory $\boldsymbol{\zeta^{0}_{t}}$, we construct a finite-horizon LQ game, which can be solved via coupled Riccati equations using the methods proposed in [36]. The linear feedback control solution of this LQ game is then used to update the agents’ policies, as suggested in [5]. This iterative process continues until the system trajectory remains unchanged at each iteration and converges to a fixed trajectory, known to be the local Nash equilibrium trajectory $\boldsymbol{\zeta^{*}_{t}}$.
@@
In an incomplete information game, both agents are learning the other agent’s parameter $\theta^{-k}$, resulting in an estimate $\hat{\theta}^{-k}_{t}$. As a result, each agent independently solves a different iLQR problem to update their policy using $\theta^{k}$ and $\hat{\theta}^{-k}_{t}$, leading to a distinct converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$ at each step. At the converged trajectories, each agent’s policy consists of a feedforward time-varying term $\alpha_{t}$ and a feedback time varying term $P_{t}$ in the form of
@@
Remark 1. The proposed framework in this paper is not limited to using iLQR for solving general-sum games. The key requirement for the success of N-PACE is to have a policy explicitly parameterized by $\theta^{k}$ and $\theta^{-k}$. In fact, alternative approaches, such as some reinforcement learning algorithms [38, 17], can also be used.
III-C Modeling the Learning Dynamics
In Sec. III-B, we discussed how an agent $k$ can update its policy at each step based on the current estimate $\hat{\theta}^{-k}_{t}$ by solving a general-sum game. However, the question of how to learn and update $\hat{\theta}^{-k}_{t}$ remains open. In this section, we propose our learning algorithm under the assumption that agent $-k$ is an expert, for the sake of clarity and brevity. We then introduce approximation methods to enhance the efficiency of this learning process for real-time applications. Finally, in Sec. III-D, we discuss how this learning approach can be extended to N-PACE for scenarios where all agents are learners.
The learning dynamics $h^{k}_{l}$. When agents can observe the states and control signals $s_{t},a^{k}_{t},a^{-k}_{t}$ at each step, they can leverage this information to construct an error term for the online learning of the parameters $\hat{\theta}^{-k}_{t}$. An agent $k$, knowing that their peer is solving an iLQR game to generate policies, can attempt to model this policy using their current estimate $\hat{\theta}^{-k}_{t}$ as $\hat{a}^{-k}_{t}=\pi^{-k}_{t}(s_{t};\hat{\theta}^{-k}_{t},\theta^{k})$ as given in (3). Consequently, they can define the following update rule as their learning dynamics $h^{k}_{l}$ to refine the estimate of $\hat{\theta}^{-k}_{t}$

### v2

@@ Abstract.
Dynamic game theory is a powerful tool in modeling multi-agent interactions and human-robot systems. In practice, since the objective functions of both agents may not be explicitly known to each other, these interactions can be modeled as incomplete-information general-sum dynamic games. Solving for equilibrium policies for such games presents a major challenge, especially if the games involve nonlinear underlying dynamics. To simplify the problem, existing work often assumes that one agent is an expert with complete information about its peer, which can lead to biased estimates and failures in coordination. To address this challenge, we propose a nonlinear peer-aware cost estimation (N-PACE) algorithm for general-sum dynamic games. In N-PACE, using iterative linear quadratic (ILQ) approximation of dynamic games, each agent explicitly models the learning dynamics of its peer agent while inferring their objective functions and updating its own control policy accordingly in real time, which leads to unbiased and fast learning of the unknown objective function of the peer agent. Additionally, we demonstrate how N-PACE enables intent communication by explicitly modeling the peer’s learning dynamics. Finally, we show how N-PACE outperforms baseline methods that disregard the learning behavior of the other agent, both analytically and using our case studies.11footnotemark: 1Codes: AAMAS2026-NPACE.
Key words and phrases:
Incomplete Information Dynamic Games, Optimal Control, Intent Inference, Learning from a Learner, ILQgames
1. Introduction
@@
General-sum dynamic games offer a powerful framework for modeling multi-agent interactions schwarting2019social; losey2018review, with applications in autonomous driving chen2021shall, assistive robots li2019differential, and social navigation fridovich2020efficient. However, the process of finding the feedback Nash equilibrium (FBNE) solution of such games, which can be achieved by solving a set of Hamilton-Jacobi-Bellman (HJB) equations bressan2010noncooperative, is inherently complex, as traditional dynamic programming approaches suffer from the curse of dimensionality powell2007approximate. This problem becomes even more challenging when agents do not know each other’s objective function and must perform inference online while making decisions, resulting in an incomplete-information differential game, a common scenario in human-robot interaction (HRI) li2019differential; nikolaidis2017game; kedia2024interact; tian2023towards and autonomous vehicle interaction amatya2022shall; wang2019enabling; xing2019driver.
@@
Figure 1. Demonstration of the core idea in N-PACE using a simulated general-sum game between two autonomous cars at an intersection. In (a), each agent assumes its peer knows its aggressiveness level, leading to biased inferences and possible collisions. In (b), N-PACE models the peer agent as a learning entity, resulting in a safe intersection crossing.
footnotetext: Extended preprint version. This manuscript is an extended version of our paper accepted to Proc. of AAMAS 2026.
@@
Existing approaches to solve such games often rely on the assumption that one agent is an expert which knows the true intent of other agents peters2024contingency; le2021lucidgames; li2024intent; schwarting2019social. Despite success of these approaches in situations where the peer agent is indeed a complete information agent, in scenarios where all agents are learning simultaneously and inferring each other’s objective functions (which is defined as their intent in this paper), these assumptions can lead to biased estimates, poor coordination, and failure to achieve each agent’s goal liu2016blame; soltanian2025pace. This is because while doing the inference, they fail to account for the fact that other agents can be learners, whose policies are conditioned on previously observed data kedia2024interact. It is even more important to consider such mutual learning in a competitive setting where an agent has no information about the others.
@@
In this paper, we highlight the importance of accounting for the learning dynamics of peer agents while inferring their objectives, addressing critical limitations in existing approaches. We achieve this goal by developing a peer-aware cost estimation algorithm for general-sum nonlinear games (N-PACE), where each agent leverages prior knowledge about its peer agent’s learning dynamics (e.g., gradient-based learner, Bayesian learner) to condition its current inference of the other’s objective on their parameter learning, thereby removing the bias in their inferences. The N-PACE algorithm employs iterative linear quadratic (ILQ) approximation of the nonlinear general-sum game fridovich2020efficient, both in policy update and intent inference phases. We also demonstrate how (ILQ) enables efficient approximation during the inference phase of our algorithm.
@@
Contributions. This paper makes the following contributions. (1) We introduce the N-PACE algorithm for mutual intent inference in multi-agent interaction with continuous action-state and continuous intent space, where all agents are making online inferences. (2) We empirically and theoretically demonstrate the advantages of N-PACE in convergence, safety, and task completion, compared to baseline methods that assume the peer agent is an expert who knows the ego agent’s true intent. (3) We demonstrate how N-PACE, along with access to the learning dynamics of the peer agent, can be leveraged for intent communication and signaling to facilitate mutual learning and adaptation.
@@
The closest works to ours are fridovich2020efficient; li2024intent; liu2016blame; soltanian2025pace, where fridovich2020efficient uses ILQGame for solving complete information multi-agent interactions. In li2024intent, the ILQ approach is employed in a general-sum game where an “expert” demonstrates its intent to a learner agent knowing their learning dynamics, but it does not consider the situation when two learning agents (no experts) try to coordinate which is a more practical setting in multi-agent interactions. Finally, while liu2016blame; soltanian2025pace study the problem of mutual learning and adaptation, the former’s problem setup is limited to a single-step optimization horizon cost optimization, and the latter considers only the linear quadratic setup. This work extends those to nonlinear general-sum games, which is a more practical multi-agent setup.
2. Related Work
@@
Solving general-sum games. Multi-agent interactions have been explored using various frameworks, including multi-agent reinforcement learning lowe2017multi; canese2021multi, adaptive control chen2019control, and game theory yang2020overview; bloembergen2015evolutionary. A large portion of these interactions can be modeled as general-sum dynamic games zhang2023approximating; zhang2024pontryagin. Traditional approaches for solving such games rely on adaptive dynamic programming to compute feedback Nash equilibria (FBNE) vamvoudakis2011multi. However, these methods are often hindered by the curse of dimensionality powell2007approximate, making them impractical for high-dimensional problems and real-time implementation in robotics. To overcome these limitations, physics-informed neural networks (PINN) have been used to solve the HJB equation more efficiently zhang2024pontryagin. Some recent work has focused on iterative linear-quadratic (ILQGames) approximations, which offer a more computationally efficient strategy for approximating the multi-agent interactions solution by finding the local Nash equilibrium of such games fridovich2020efficient. In this paper, we employ ILQGame for both the learning and control phases to find the approximated Nash equilibrium policies.
@@
Incomplete information general-sum games. Solving incomplete information general-sum games is significantly more challenging than their complete-information counterparts aumann1995repeated; cardaliaguet2012games. These games often require simplifications, such as modeling one agent (the ego agent) as a learner while assuming others are fully informed peters2024contingency; le2021lucidgames; laine2021multi, treating the problem as an adaptive control problem li2019differential, discretizing the intent space or relying on offline datasets for intent inference and belief updates hu2023emergent; mehr2023maximum; chen2021shall. Our work is motivated by the lack of studies on incomplete information dynamic games where all agents are learners jacq2019learning. We will highlight the potential issues caused by these simplifications, such as modeling other agents as fully informed.
@@
Learning from a learner. Online inference of an expert agent’s objective function has been extensively studied through inverse optimal control and inverse reinforcement learning molloy2020online; rhinehart2017first; self2022model. While inverse non-cooperative games have been explored both offline and online molloy2022inverse; li2023cost, existing works focus on learning from an expert system. Some works have considered opponent shaping and learning from a learner agent by incorporating their learning dynamics into offline multi-agent policy training approaches foerster2017learning; jacq2019learning, but they are inherently offline multi-agent policy training algorithms and not primarily designed for real-time intent inference and objective learning. Moreover, they are not built on game-theoretic setups where the future action of one agent is conditioned on others. But motivated by these learning from a learner approaches and recent efforts in modeling human learning dynamics in human-robot interactions tian2023towards; kedia2024interact, we propose a framework for mutual intent inference in systems where all agents are learners.
3. Problem Formulation
@@
Notations and Assumptions. We focus on a two-player setting with agents $i$ and $j$. When referring to an arbitrary agent, we use the index $k$ and denote their counterpart as k-k. We consider the following system dynamics
@@
The Feedback Nash Equilibrium (FBNE) bacsar1998dynamic is a solution concept for the above minimization problem, constrained by (1), where neither agent has an incentive to deviate from their policy, as doing so would increase their cumulative cost. In an FBNE, each agent’s policy is a function of the current state, ensuring that at every time step, the chosen control actions are optimal given the strategies of the other agents. We represent these policies using $\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$ such that $a^{k}_{t}=\pi^{k}_{t}(s_{t};\theta^{i},\theta^{j})$, emphasizing that they depend on both $\theta^{i}$ and $\theta^{j}$, which is crucial for further formulation in our work.
@@
The incomplete information game. The previously discussed FBNE solution was derived under the assumption that both agents have full knowledge of $\theta^{i}$ and $\theta^{j}$. However, in incomplete information games, $\theta^{j}$ may be unknown to agent $i$, and $\theta^{i}$ may be unknown to agent $j$. As a result, at each time step, they must rely on estimations $\hat{\theta}^{j}_{t}$ and $\hat{\theta}^{i}_{t}$, respectively, and update these estimates over time. This leads to policies of the form $\pi^{i}_{t}(s_{t};\theta^{i},\hat{\theta}^{j}_{t})$ and $\pi^{j}_{t}(s_{t};\hat{\theta}^{i}_{t},\theta^{j})$, which are not necessarily the same as FBNE strategies. To address this problem, the next section introduces how we use the ILQ method to obtain the policy $\pi^{k}_{t}(s_{t};\theta^{k},\hat{\theta}^{-k}_{t})$ for each agent $k$, based on their current estimate of the other agent’s parameter $\hat{\theta}^{-k}_{t}$. We then propose appropriate update rules for these estimates in Sec. 3.2.
3.1. ILQGame for Multi-Agent Interactions
@@
The iterative linear quadratic regulator (iLQR) approach for locally optimal feedback control of nonlinear systems was first introduced in li2004iterative. More recent work, such as fridovich2020efficient, has demonstrated its efficiency in computing local Nash equilibrium solutions for general-sum dynamic games in real time. In our incomplete information game, where we need policies $\pi^{k}_{t}$ to be computed as functions of the estimated parameters $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$ in real time, we leverage the proposed ILQGame algorithm in fridovich2020efficient to solve the game for each agent $k$ at every time step. The main reasons for using ILQGame in this work are its efficiency in real-time computations and, more importantly, its capability to parameterize the obtained policies as functions of $\hat{\theta}^{i}_{t}$ and $\hat{\theta}^{j}_{t}$, which is crucial for having a differentiable layer in N-PACE as shown in Sections 3.2 and 3.3.
@@
In the ILQGame approach, we begin with an initial admissible policy for each agent. In the forward pass, by applying these policies to the dynamics (1), we obtain a trajectory $\boldsymbol{\zeta^{0}_{t}}=\{(s^{0}_{\tau},a^{i0}_{\tau},a^{j0}_{\tau})\mid\tau\in\{t,t+1,\dots,T\}\}$. Next, by linearizing the dynamics $f_{t}$ and quadratizing the running costs $g^{k}_{t}$ in the neighborhood of the trajectory $\boldsymbol{\zeta^{0}_{t}}$, we construct a finite-horizon LQ game, which can be solved via coupled Riccati equations using the methods proposed in bacsar1998dynamic. The linear feedback control solution of this LQ game is then used to update the agents’ policies, as suggested in fridovich2020efficient. This iterative process continues until the system trajectory remains unchanged at each iteration and converges to a fixed trajectory, known to be the local Nash equilibrium trajectory $\boldsymbol{\zeta^{*}_{t}}$.
@@
In an incomplete information game, each agent learns the other’s parameter $\theta^{-k}$, resulting in an estimate $\hat{\theta}^{-k}_{t}$. As a result, each agent independently solves a different ILQ problem to update their policy using $\theta^{k}$ and $\hat{\theta}^{-k}_{t}$, leading to a distinct converged trajectory $\zeta^{*k}_{t}(\theta^{k},\hat{\theta}^{-k}_{t})$ at each step. At the converged trajectories, each agent’s policy consists of a feedforward time-varying term $\alpha_{t}$ and a feedback time varying term $P_{t}$ in the form of
@@
Remark 0.
(Extension to Other Methods) The core idea in this paper is not limited to using ILQ for solving general-sum games. The key requirement for the success of N-PACE is to have a policy explicitly parameterized by $\theta^{k}$ and $\theta^{-k}$. In fact, alternative approaches, such as some reinforcement learning algorithms williams1992simple; lowe2017multi, can also borrow the N-PACE core idea as long as they can generate policies parametrized by intent parameters.
3.2. Modeling the Learning Dynamics
In Sec. 3.1, we discussed how an agent $k$ can update its policy at each step based on the current estimate $\hat{\theta}^{-k}_{t}$ to solve a general-sum game. However, the question of how to learn and update $\hat{\theta}^{-k}_{t}$ remains open. In this section, we propose a learning algorithm under the assumption that agent k-k is an expert, for the sake of clarity and brevity. We then introduce approximation methods to enhance the efficiency of this learning process for real-time applications. Finally, in Sec. 3.3, we discuss how this learning approach can be extended to N-PACE for scenarios where all agents are learners.
The learning dynamics $h^{k}_{l}$. When agents can observe the states and control signals $s_{t},a^{k}_{t},a^{-k}_{t}$ at each step, they can leverage this information to construct an error term for the online learning of the parameters $\hat{\theta}^{-k}_{t}$. An agent $k$, knowing that their peer is solving an ILQGame to generate policies, can attempt to model this policy using their current estimate $\hat{\theta}^{-k}_{t}$ as $\hat{a}^{-k}_{t}=\pi^{-k}_{t}(s_{t};\hat{\theta}^{-k}_{t},\theta^{k})$ as given in (3). Consequently, they can define the following update rule as their learning dynamics $h^{k}_{l}$ to refine the estimate of $\hat{\theta}^{-k}_{t}$